14. A dime has a diameter of 1.8000 cm at 20C. The dime is made of mostly zinc, for which the 

coefficient of linear expansion is 2.6 10 −5C−1.  

 

(a) What would its diameter be on a hot day at 48.0C?  

(b) What is its diameter on a cold night at −53.0C? 



1
Expert's answer
2021-07-13T10:19:51-0400

Solution.

"D_0=1.8000cm;"

"T_0=20^oC;"

"\alpha=2.61\sdot10^{-5}C^{-1};"

"a)T=48.0^oC;"

"\Delta D=\alpha D_0\Delta T;"

"\Delta D=2.61\sdot10^{-5}\sdot1.8000\sdot(48.0-20)=0.0013cm;"

"D=D_0+\Delta D;"

"D=1.8000+0.0013=1.8013cm;"

"b) T=-53.0^oC;"

"\Delta D=\alpha D_0\Delta T;"

"\Delta D=2.61\sdot10^{-5}\sdot1.8000\sdot(-53.0-20)=-0.0034cm;"

"D=D_0+\Delta D;"

"D=1.8000-0.0034=1.7966cm;"

Answer: "a)D=1.8013cm;"

"b)D=1.7966cm."